TMS Meeting Minutes, April 8, 2021

Attendees: Andrew, Breann, Heejoo, John, Matthieu, Olga, Trenton, Yang (Zoom Meeting)


April 13-15, Heejoo Visiting LBT Tuesday to Thursday
  • Planning to complete TMS install on SX side
  • Mike Gardiner has overlapping time with Heejoo; planning to work together, Mike will be bringing the laser tracker
  • Heejoo is still waiting for the delivery of the custom collimators from Thorlabs; they are supposed to arrive next week, after the trip
  • Heejoo planning to perform some basic collimator tests off the mountain and then ask Mike or Leroy to test on the mountain once he sends one up to the mountain
  • Heejoo will open a Zoom session for the laser tracker survey for all interested parties
  • Andrew requests a 3rd measurement with the solid RR at 20 m distance
    • Will arrange for the RR to be shipped to GMTO in Pasadena after that
  • Please send Heejoo any further requests you have for him before he heads up to the mountain

Passive Data, 4/3 & 4/4 UT
  • Data up on Google Drive, including mirror diameter
  • Andrew hoping to fit thermal to this data, very interested in correlation of glasstherm with focus
    • Andrew needs every bit of thermal information on the mirror possible to extract from DMS
    • Note, thermistors are very accurate and have minimal noise (only 6 thermistors on each side)
    • The thermocouples have very poor accuracy (there are 72 of them spread throughout each mirror); they have gotten old and have all sorts of problems
    • John notes that there is historical thermocouple data in telemetry (back to 2014) which is much more reliable
    • Heejoo will download these thermal datasets for Andrew
    • Andrew optimistic that we can reduce the thermal focus below 1000 nm's
  • Airpath-corrected diameter, integrated, should be correlated to the mean temperature of the glass

Matthieu Updated TMS Laser Warning Email
  • The email now includes the last received voltage from the iMonit laser power meter
  • If the readout is 0 V, we are safe and now action is needed
  • If not 0 V, we may have a problem

Etalon Software Crash
  • Doesn't seem to happen during TMS operation, only following
  • Have yet to receive a response from Etalon about this matter
  • Heejoo planning to send a follow up to touch base

John: 37.9 um of mirror Z motion corresponds to 1000 nm Z4

Yang's Software Proposal
  • Yang has prepared a draft of the software update proposals to send to Etalon
  • If you have any input, please include it in the next few days for inclusion in the email

Double-Sided TMS Code
  • At minimum, you could launch two instances of the TMS software, one for each side, computed separately
  • Andrew clarifies that he would like to see measurements on two mirrors coming in at the same time, and we need to be able to separate them out for separate inverse kinematics calculations
    • Want a single measurement request (in one instance of the software), with a single vector of lengths from both sides, then process them separately, separating out lenghts according to the appropriate channel lengths
    • Notes that this could result in the maximum cadence of measurement
    • There is also the potential issue of measurements colliding from the two sides; but the Multiline software should handle this fine by simply queing the measure commands
  • Yang plans to think on how this should be restructured to accomodate these changes and work with John's program (John's program is already configured to work for both sides)
  • Also planning to enable the ability to switch sides, or run both sides at once

Summer Break Plans
  • John states that we should assume there will be a normal summer shutdown, predicated on having no new wave of COVID in AZ
  • For mirror coating or cleaning, all of the TMS hardware needs to be removed
  • Heejoo planning to write some instructions on how to remove the collimators and reinstall them
  • Andrew's request: PLEASE re-number the collimators to match up position channel number with the Etalon channel number
  • Heejoo would like to re-number the channels to simplify everything and reduce the chance for mistakes
  • Andrew suggests keeping the fibers connected to the collimators and just rearranging them, limiting risk of accidentally defocusing the collimator-fiber system
  • There are three broken 70 m fibers, Heejoo planning to replace them this summer
  • We are not currently using the MPO casette, and it would be really nice to have that working; plan to work on this over the summer (may need more collimators to look at LBTI rim)
  • We have some time to prepare this plan, so please provide input on any action items you'd like to see or any potential TMS upgrades
    • Andrew: Set up a different collimator arrangement with a flat mirror/90 deg prism to orient the collimators tangential to the beam, eliminating fiber rubbing on the edge of the mirror hub

-- TrentonBrendel - 08 Apr 2021
Topic revision: r2 - 15 Apr 2021, HeejoChoi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback